(a) Calculate the angular momentum of Earth that arises from its spinning motion on its axis, treating Earth as a uniform solid sphere. J · s

(b) Calculate the angular momentum of Earth that arises from its orbital motion about the Sun, treating Earth as a point particle. J · s
Answer #1

a)

The angular momentum of Earth can be obtained by

L=1w

Where

I = moment of inertia of a sphere (as Earth is considered a sphere) = 2/5 mr2

\omega = angular velocity = 27/1day = 27/24 * 3600) ( as 1 day as 24 hours and each hour has 3600 s)

So,

L = 2/5 mr2 x 2r (24 * 3600)

m = mass of earth = 5.97 x 1024 kg

r = radius of earth = 6.38 x 106 m

So, L = 2/5 x (  5.97 x 1024 x (6.38 x 106)2 ) x 2r (24 * 3600) Js

So, L = 7.07 x 1033 Js-----------------(Answer)

b)

We will use the same expression for angular momentum but in this case

I = moment of inertia = mr2

\omega = angular velocity = 27/365days = 27/365 * 24 * 3600) ( as revolution time is 365 days and 1 day has 24 hours and each hour has 3600 s)

So,

L = mr2 x 2 (365 + 24 + 3600)

m = mass of earth = 5.97 x 1024 kg

r = radius of orbit = 1.5 x 1011 m

So, L = (  5.97 x 1024 x (6.38 x 106)2 ) x 2 (365 + 24 + 3600) Js

So, L = 2.676 x 1040 Js----------------(Answer)